OpenStudy (anonymous):

Can you check my work? √(x+2) + 4 = x. Does x = 7?

jimthompson5910 (jim_thompson5910):

Let's check that potential answer √(x+2) + 4 = x √(7+2) + 4 = 7 .. replace each x with 7 √(9) + 4 = 7 3 + 4 = 7 7 = 7 so the answer x = 7 is confirmed

OpenStudy (loser66):

@jim_thompson5910 when solving as quadratic, I cannot get that answer, why? Please, explain me

OpenStudy (anonymous):

@Loser66 did you square correctly? did you get two answers?

OpenStudy (loser66):

oh yea !@satellite73 my bad. I am sorry. XD

OpenStudy (goformit100):

"Welcome to OpenStudy. I can guide regarding this useful site; ask your doubts from me, for it you can message me. Please use the chat for off topic questions. And remember to give the helper a medal, by clicking on "Best Answer". We follow a code of conduct, ( http://openstudy.com/code-of-conduct ). Please take a moment to read it."